# Scaling and Load-Balancing Equi-Joins

### v2

@@ 8.2.5 Performance on Small-Large Outer-Joins
8.3 Experiments on Real Data
Figure 15: The runtime of the equi-join algorithms on $1000$ executors on real data.
@@ 8.3 Experiments on Real Data
Table 4: The frequencies of the hottest $10$ keys used in real-data experiments.
The second experiment was a join between two relations. The first relation is the one described above. The second relation represented data on a single day, October $14^{th}$, 2020. The size of the second relation was $0.57$ TB. The runtime of the equi-join algorithms on $1000$ executors is shown in Fig. 15. All the algorithms, were able to finish successfully, except for Broadcast-Join since it could not fit the smaller relation in the memory of the executors. Thanks to organizing the equi-join execution in multiple stages, AM-Join and Tree-Join were able to execute an order of magnitude faster than all the other algorithms. Their basic counterparts however, executed significantly slower due to the load imbalance of building the initial index and splitting the hot keys in the first iteration. ExpVar-Join executed faster than the basic counterparts but slower than AM-Join and Tree-Join. The large number of hot keys in the real relations slowed ExpVar-Join since the complexity of its pre-join step grows with the number of hot keys.
9 Conclusion
This paper proposes Adaptive-Multistage-Join (AM-Join) a fast, efficient and scalable equi-join algorithm that is built using the basic MapReduce primitives, and is hence deployable in any distributed shared-nothing architecture. The paper started by proposing Tree-Join, a novel algorithm that organizes the equi-join execution in multiple stages. Tree-Join attains scalability by distributing the load of joining a key that is hot in both relations throughout the join execution. Such keys are the scalability bottleneck of most of the state-of-the-art distributed algorithms. AM-Join utilizes Tree-Join for load balancing, high resource utilization, and scalability. Moreover, AM-Join utilizes Broadcast-Joins that reduce the network load when joining keys that are hot in only one relation. By utilizing Tree-Join and Broadcast-Join, AM-Join achieves speed, efficiency, and scalability. AM-Join extends to all the outer-joins elegantly without record deduplication or custom table partitioning, unlike the state-of-the-art industry-scale algorithms [21]. For the fastest execution of AM-Join outer-joins, the paper proposed Index-Broadcast-Join (IB-Join) family that improves on the state-of-the-art Small-Large outer-join algorithms [91, 27]. All the proposed algorithms use the basic MapReduce primitives only, and hence can be adopted on any shared-nothing architecture.
@@ 9 Conclusion
All the theoretical claims have been verified using extensive evaluation. Our evaluation highlights the improved performance and scalability of AM-Join when applied to the general equi-joins. When compared to the sate-of-the-art algorithms [21, 41], AM-Join executed comparably fast on weakly-skewed synthetic tables and can join more-skewed or orders-of-magnitude bigger tables, including our real-data tables. These advantages are even more pronounced when applying the join algorithms to natural self-joins. The proposed IB-Join outer-join algorithm executed much faster than the state-of-the-art algorithms in [91, 27].
Our future directions focus on optimizing AM-Join for the general shared-nothing architecture that supports multicasting data, for NUMA machines connected by a high-bandwidth network, and learning from the RDMA and the work-stealing enhancements of [71]. Moreover, we plan to explore using Radix join [59, 13] that it is only bound by the memory bandwidth as the Shuffle-Join.
10 Acknowledgments

### v3

@@ 8.2.5. Performance on Small-Large Outer-Joins
8.3. Experiments on Real Data
Figure 15. The runtime of the equi-join algorithms on $1000$ executors on real data.
@@ 8.3. Experiments on Real Data
Table 4. The frequencies of the hottest $10$ keys used in real-data experiments.
@@
The second experiment was a join between two relations. The first relation is the one described above. The second relation represented data on a single day, October $14^{th}$, 2020. The size of the second relation was $0.57$ TB. The runtime of the equi-join algorithms on $1000$ executors is shown in Fig. 15. All the algorithms were able to finish successfully, except for Broadcast-Join since it could not fit the smaller relation in the memory of the executors. Thanks to organizing the equi-join execution in multiple stages, AM-Join and Tree-Join were able to execute an order of magnitude faster than all the other algorithms. Their basic counterparts however, executed significantly slower due to the load imbalance of building the initial index and splitting the hot keys in the first iteration. ExpVar-Join executed faster than the basic counterparts but slower than AM-Join and Tree-Join. The large number of hot keys in the real relations slowed ExpVar-Join since the complexity of its pre-join step grows with the number of hot keys.
9. Conclusion
This paper proposes Adaptive-Multistage-Join (AM-Join) a fast, efficient and scalable equi-join algorithm that is built using the basic MapReduce primitives, and is hence deployable in any distributed shared-nothing architecture. The paper started by proposing Tree-Join, a novel algorithm that organizes the equi-join execution in multiple stages. Tree-Join attains scalability by distributing the load of joining a key that is hot in both relations throughout the join execution. Such keys are the scalability bottleneck of most of the state-of-the-art distributed algorithms. AM-Join utilizes Tree-Join for load balancing, high resource utilization, and scalability. Moreover, AM-Join utilizes Broadcast-Joins that reduce the network load when joining keys that are hot in only one relation. By utilizing Tree-Join and Broadcast-Join, AM-Join achieves speed, efficiency, and scalability. AM-Join extends to all the outer-joins elegantly without record deduplication or custom table partitioning, unlike the state-of-the-art industry-scale algorithms (Bruno et al., 2014). For the fastest execution of AM-Join outer-joins, the paper proposed Index-Broadcast-Join (IB-Join) family that improves on the state-of-the-art Small-Large outer-join algorithms (Xu and Kostamaa, 2010; Cheng et al., 2017). All the proposed algorithms use the basic MapReduce primitives only, and hence can be adopted on any shared-nothing architecture.
@@ 9. Conclusion
All the theoretical claims have been verified using extensive evaluation. Our evaluation highlights the improved performance and scalability of AM-Join when applied to the general equi-joins. When compared to the sate-of-the-art algorithms (Bruno et al., 2014; Gavagsaz et al., 2019), AM-Join executed comparably fast on weakly-skewed synthetic tables and can join more-skewed or orders-of-magnitude bigger tables, including our real-data tables. These advantages are even more pronounced when applying the join algorithms to natural self-joins. The proposed IB-Join outer-join algorithm executed much faster than the state-of-the-art algorithms in (Xu and Kostamaa, 2010; Cheng et al., 2017).
Our future directions focus on optimizing AM-Join for the general shared-nothing architecture that supports multicasting data, for NUMA machines connected by a high-bandwidth network, and learning from the RDMA and the work-stealing enhancements of (Rödiger et al., 2016). Moreover, we plan to explore using Radix join (Manegold et al., 2002; Barthels et al., 2015) that it is only bound by the memory bandwidth as the Shuffle-Join.
Acknowledgements.